### v1

# Thermodynamics Reconstructed from Information Science: Theoretical Core

Abstract
We present a minimal information-theoretic/axiomatic framework that reconstructs thermodynamic structure from a pair of primitive data: a measurement (coarse-graining) map $M$ and a reference measure (prior) $\tau$. For each cell $C_{y}=M^{-1}(y)$ in the state space, we introduce an information volume $\tau(C_{y})$ and its log-volume, and incorporate it—on the same footing as energy and particle number—as an expectation constraint in a minimum relative-entropy (Min-KL) problem. Temperature, chemical potential, and pressure are then defined primarily as dual variables of a relative-entropy functional, yielding a Legendre structure from which a first-law-type differential relation for internal energy follows. In particular, pressure is reinterpreted as the variable conjugate to an information volume rather than a geometric volume.
For nonequilibrium processes, we define the path-space relative entropy $\Sigma_{0,T}$ between forward and time-reversed path measures as entropy production at the path level, and decompose $\Sigma_{0,T}$ into the sum of the system entropy change and an environment entropy change using observational entropy $S_{M,\tau}$. With temperature introduced as the static Min-KL dual variable, we define heat as $Q:=T\,\Delta S_{\mathrm{env}}$, thereby reconstructing the conventional heat quantity as the representation $Q=T(\Sigma_{0,T}-\Delta S_{\mathrm{sys}})$ without model-dependent assumptions such as local detailed balance.
We further compare path-space relative entropies of a joint system and a subsystem, and show that their difference is expressed as an expectation of the KL divergence between conditional path measures, proving nonnegativity. This unifies Ito–Sagawa-type information terms and “hidden dissipation” as an information-geometric gap induced by projection from the joint description to a marginal one. On the static side, for additive Hamiltonians we formulate the exact decomposition of the joint nonequilibrium free energy into subsystem free energies plus a correlation free-energy term proportional to the mutual information $I(X;Y)$, understood as a KL-geometric decomposition.
Finally, we treat the choice of $(M,\tau)$ as a gauge (convention), distinguish gauge-invariant quantities (e.g. path relative entropy and free-energy differences) from gauge-dependent bookkeeping terms, and present a two-state Markov-chain toy model where $\Sigma_{0,T}$, observational dissipation, hidden dissipation, and heat $Q$ are computed explicitly.
1 Introduction
Thermodynamic quantities (temperature, chemical potential, pressure, heat, work, etc.) have traditionally been defined either (i) through equilibrium statistical mechanics, via partition functions and thermodynamic potentials, or (ii) on specific dynamical models describing energy exchange with a heat bath (e.g. Markov processes satisfying local detailed balance (LDB)). From the viewpoint of information theory, by contrast, entropy and free energy are naturally characterized by the Kullback–Leibler (KL) divergence, and fluctuation theorems and the second law are often expressed as time-reversal asymmetry of path measures [9, 10, 11, 13].
The aim of this paper is to push this perspective one step further:
  1. (1)
    We introduce an information-volume variable derived from primitive data $(M,\tau)$—a measurement map $M$ and a reference measure $\tau$—and define pressure as a primary dual variable, symmetric to temperature and chemical potential, through an expectation constraint on the log-volume.
  2. (2)
    We introduce dissipation as the path-space KL divergence $\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)$ in a primary manner, and define the environment entropy change $\Delta S_{\mathrm{env}}$ as a representation theorem when combined with observational entropy $S_{M,\tau}$.
  3. (3)
    We integrate these static and dynamic constructions under a single reference measure $\tau$, and organize their relations to existing information thermodynamics (observational entropy, mutual information, hidden dissipation, etc.) in an axiomatic way.
We present static Min-KL inference, dynamic path-KL dissipation, and the resulting information-theoretic representation of heat in as model-independent a form as possible.
2 Proposed model and theoretical framework: primitive data, information volume, and observational entropy
2.1 Primitive data
We fix the following three items as primitive data:
  1. 1.
    State space: a measurable space $(\Omega,\mathcal{F})$.
  2. 2.
    Measurement (coarse-graining): a measurable map $M:\Omega\to\mathcal{Y}$ (including finite resolution, partitions, and measurement).
  3. 3.
    Reference (coding convention / prior): a reference measure $\tau$ (assumed $\sigma$-finite).
The measurement map $M$ specifies which degrees of freedom are distinguished by observation. It associates a microstate $x\in\Omega$ with an observed outcome $y=M(x)$; thus $M$ represents, in a unified manner, “cell partitions of phase space,” “instrument resolution,” and “the choice of coarse-graining.” For example, if in a one-dimensional particle system the position is only measured on a mesh of size $\Delta x$, then $M$ may be regarded as mapping a continuous coordinate $x$ to a discrete label $y\in\mathbb{Z}$.
The reference measure $\tau$ encodes a prior (coding convention): “which states are regarded as uniform” at the descriptive level. For continuous systems, the Liouville or Lebesgue measure is a natural candidate; for discrete systems one often chooses the uniform distribution or an equilibrium distribution. In this paper, we introduce $\tau$ primarily as an information-theoretic reference measure, and assign physical meaning (equilibrium vs. nonequilibrium, etc.) as needed later.
Let the cell (equivalence class) induced by $M$ be
$C_{y}\;:=\;M^{-1}(y).$ (1)
The prior $\tau$ then induces a “volume” of each cell,
$v(y)\;:=\;\tau(C_{y}),$ (2)
which is not a geometric volume but a measure determined by the observation and coding convention. We call $v(y)$ the information volume.
Moreover, for each microstate $x$ we define the associated log-volume
$\ell_{V}(x)\;:=\;\log\tau\!\bigl(C_{M(x)}\bigr)\;=\;\log v\!\bigl(M(x)\bigr).$ (3)
This is the “information-volume” variable used throughout the paper.
2.2 Observational entropy
Given a microdistribution $p$ (with $p\ll\tau$), define the macrodistribution
$p_{M}(y)\;:=\;p(C_{y}).$ (4)
A representative form of observational entropy is
$S_{M,\tau}(p)\;:=\;-\sum_{y\in\mathcal{Y}}p_{M}(y)\,\log\!\left(\frac{p_{M}(y)}{\tau(C_{y})}\right)\;=\;-\sum_{y}p_{M}(y)\log p_{M}(y)\;+\;\sum_{y}p_{M}(y)\log\tau(C_{y}),$ (5)
for finite $\mathcal{Y}$. The framework and properties of observational entropy, including extensions to quantum settings and general priors, have been systematically developed by Safránek and collaborators [3, 4, 5, 6, 7]. In this paper, we take $S_{M,\tau}$ as the primary definition of “system entropy” in both static and dynamic discussions.
2.3 Gauge freedom in the choice of $M$ and $\tau$
The choice of the measurement map $M$ and the reference measure $\tau$ essentially fixes a gauge of description (a convention). We summarize which quantities depend on this choice and which are invariant.
(i) Refinement/coarse-graining of the measurement map
We say that $M_{1}:\Omega\to\mathcal{Y}_{1}$ is a refinement of $M_{2}:\Omega\to\mathcal{Y}_{2}$ if each equivalence class $C_{y}^{(2)}:=M_{2}^{-1}(y)$ is further partitioned into cells $C_{y^{\prime}}^{(1)}:=M_{1}^{-1}(y^{\prime})$. Then $M_{1}$ corresponds to a finer observation than $M_{2}$.
According to Safránek et al. [5], under suitable conditions one has
$S_{M_{1},\tau}(p)\;\leq\;S_{M_{2},\tau}(p).$
Thus observational entropy can increase under coarse observation. Consequently, the system entropy change
$\Delta S_{\mathrm{sys}}=S_{M,\tau}(p_{T})-S_{M,\tau}(p_{0})$
depends on the choice of $M$. Likewise, the environment entropy change $\Delta S_{\mathrm{env}}$ introduced in Sec. 5 also changes with $M$.
In contrast, the path-KL dissipation
$\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)$
is determined solely by the forward and reverse path measures $\mathbb{P},\mathbb{P}^{\mathrm{R}}$ and does not depend on $M$. In this sense, $\Sigma_{0,T}$ is gauge-invariant under changes of observation, while $\Delta S_{\mathrm{sys}}$ and $\Delta S_{\mathrm{env}}$ are gauge-dependent bookkeeping terms that allocate dissipation between “system” and “environment.”
(ii) Changing the reference measure and invariant content
Next, consider changing the reference measure from $\tau$ to $\tau^{\prime}$. Assume $\tau$ and $\tau^{\prime}$ are mutually absolutely continuous, so that for some integrable $\varphi$,
$\mathrm{d}\tau^{\prime}(x)=e^{-\varphi(x)}\,\,\mathrm{d}\tau(x).$
Then for any distribution $p$,
$D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau^{\prime}\right)=D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau\right)+\int\varphi(x)\,p(\,\mathrm{d}x)+\text{(constant)}.$
Hence the Min-KL structure is preserved, while the numerical value of the optimum $D_{\mathrm{KL}}\!\left(p^{\ast}\,\middle\|\,\tau\right)$ generally depends on $\tau$.
In this paper we fix $\tau$ as a coding convention specifying “uniformity” and define $S_{M,\tau}$, $\ell_{V}$, pressure $P$, etc. consistently under that choice. Switching to another $\tau^{\prime}$ corresponds to a gauge change of the theory; physical content can be extracted by focusing on quantities invariant under this change (e.g. entropy differences between two states, and the path-KL dissipation $\Sigma_{0,T}$).
Remark 1.
In the decomposition
$\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\Delta S_{\mathrm{env}},$
both $\Delta S_{\mathrm{sys}}$ and $\Delta S_{\mathrm{env}}$ depend on the choice of $M$ and $\tau$. However, $\Sigma_{0,T}$ itself is uniquely determined once the forward and reverse path measures are given and is invariant under gauge changes of $(M,\tau)$. Thus our framework separates the total dissipation from its allocation between system and environment.
3 Static theory: Min-KL inference and dual variables
We define thermodynamic states not by dynamical assumptions but as outcomes of an information-update rule. The basic principle is minimum cross-entropy (minimum KL) with respect to the reference measure $\tau$, justified information-theoretically by Jaynes’ MaxEnt interpretation [1] and the Shore–Johnson axiomatization [2].
3.1 Variational problem (state as a definition)
As constraints, we specify expectations of energy $E(x)$, particle number $N(x)$, and the information-volume code length $\ell_{V}(x)$:
$\mathbb{E}_{p}[E]=U,\qquad\mathbb{E}_{p}[N]=\bar{N},\qquad\mathbb{E}_{p}[\ell_{V}]=\bar{v}.$ (6)
We then define the state $p^{\ast}$ by
[…]
Here the KL divergence (using the Radon–Nikodym density $r=\,\mathrm{d}p/\,\mathrm{d}\tau$) is
$D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau\right)\;:=\;\int_{\Omega}\log r(x)\,p(\,\mathrm{d}x)\;=\;\int_{\Omega}r(x)\log r(x)\,\tau(\,\mathrm{d}x),\qquad r=\frac{\,\mathrm{d}p}{\,\mathrm{d}\tau}.$ (8)
3.2 Exponential-family solution and dual variables
Under standard convexity and regularity conditions, the solution belongs to an exponential family:
$p^{\ast}(\,\mathrm{d}x)\;=\;\frac{1}{Z(\beta,\alpha,\pi)}\,\exp\!\Bigl(-\beta E(x)-\alpha N(x)-\pi\ell_{V}(x)\Bigr)\,\tau(\,\mathrm{d}x),$ (9)
where $Z$ is the normalization constant. The multipliers $\beta,\alpha,\pi$ are Lagrange multipliers (shadow prices / dual variables); general conjugacy relations for continuous MaxEnt are discussed in [8].
3.3 Definitions of temperature, chemical potential, and pressure
Our policy is to define temperature, chemical potential, and pressure primarily as dual variables, rather than introducing them a posteriori (we set $k_{\mathrm{B}}=1$):
$\frac{1}{T}\;:=\;\beta,\qquad-\frac{\mu}{T}\;:=\;\alpha,\qquad\frac{P}{T}\;:=\;\pi.$ (10)
$\mu$: shadow price of the particle-number constraint $\mathbb{E}[N]=\bar{N}$.
$P$: shadow price of the information-volume constraint $\mathbb{E}[\ell_{V}]=\bar{v}$.
In this sense, pressure $P$ is defined information-theoretically as the increase rate of KL with respect to the log-volume constraint, without relying on a geometric notion of volume.
3.4 Dual potential and a first-law-type relation
Define the log-partition function
$\Psi(\beta,\alpha,\pi)\;:=\;\log\int_{\Omega}\exp\!\bigl(-\beta E(x)-\alpha N(x)-\pi\ell_{V}(x)\bigr)\,\tau(\,\mathrm{d}x)\;=\;\log Z(\beta,\alpha,\pi).$ (11)
Under regularity,
$U=-\partial_{\beta}\Psi,\qquad\bar{N}=-\partial_{\alpha}\Psi,\qquad\bar{v}=-\partial_{\pi}\Psi.$ (12)
Let the optimal value (maximum relative entropy) be
$S^{\ast}(U,\bar{N},\bar{v})\;:=\;-D_{\mathrm{KL}}\!\left(p^{\ast}\,\middle\|\,\tau\right).$ (13)
Convex duality yields the differential relation
$\mathrm{d}S^{\ast}=\beta\,\,\mathrm{d}U+\alpha\,\,\mathrm{d}\bar{N}+\pi\,\,\mathrm{d}\bar{v}.$ (14)
Substituting (10) gives
$\mathrm{d}U=T\,\,\mathrm{d}S^{\ast}-P\,\,\mathrm{d}\bar{v}+\mu\,\,\mathrm{d}\bar{N}.$ (15)
Proposition 2 (Information-theoretic derivation of a first-law-type relation).
Assume the Min-KL problem (7) has a unique solution $p^{\ast}$, and the corresponding log-partition function $\Psi(\beta,\alpha,\pi)$ is a $C^{2}$ convex function. Let the optimal value $S^{\ast}(U,\bar{N},\bar{v})$ be defined by (13). Then the total differential in $(U,\bar{N},\bar{v})$ satisfies (14). Moreover, if intensive variables are defined by (10), the total differential of internal energy $U$ satisfies (15).
Proof.
By convex duality, $S^{\ast}$ is given by the Legendre transform
$S^{\ast}(U,\bar{N},\bar{v})=\inf_{\beta,\alpha,\pi}\bigl\{\beta U+\alpha\bar{N}+\pi\bar{v}-\Psi(\beta,\alpha,\pi)\bigr\}.$
Under regularity, the optimal $(\beta,\alpha,\pi)$ is unique, and equivalently to (12) one has $\partial_{U}S^{\ast}=\beta$, $\partial_{\bar{N}}S^{\ast}=\alpha$, $\partial_{\bar{v}}S^{\ast}=\pi$. Hence the total differential is (14).
Substituting (10) yields
$\mathrm{d}S^{\ast}=\frac{1}{T}\,\,\mathrm{d}U-\frac{\mu}{T}\,\,\mathrm{d}\bar{N}+\frac{P}{T}\,\,\mathrm{d}\bar{v}.$
Multiplying by $T$ and solving for $\mathrm{d}U$ gives (15). ∎
4 Nonequilibrium extension: path-space KL and hidden dissipation
We extend the preceding information-theoretic framework to nonequilibrium processes with time evolution. Irreversibility (entropy production, dissipation) is naturally defined as time-reversal asymmetry of stochastic processes; the most general coordinate-free formulation is given by the KL divergence between the forward and time-reversed path measures [9, 10, 11].
4.1 Definition of path dissipation (entropy production)
Consider a stochastic process defined on a time interval $[0,T]$. Let $\mathbb{P}$ be the forward path measure on path space, and let $\mathbb{P}^{\mathrm{R}}$ be the path measure of the time-reversed process induced by the time-reversal map. We define
$\Sigma_{0,T}\;:=\;D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\;\geq 0$ (16)
as entropy production (dissipation) on $[0,T]$. This definition does not invoke any model-specific assumptions such as bath temperature or LDB.
4.2 Lower bound under observation (coarse-graining) and hidden dissipation
We represent observation as a path map (channel)
$\Pi:\ \omega=(x_{0:T})\longmapsto\gamma=(y_{0:T}),\qquad y_{t}=M(x_{t}),$ (17)
where $M$ is the measurement map in (1). Write the observed path measures as $\Pi\mathbb{P}$ and $\Pi\mathbb{P}^{\mathrm{R}}$. Define observational dissipation (entropy production at the observation level) by
$\Sigma_{\mathrm{obs}}\;:=\;D_{\mathrm{KL}}\!\left(\Pi\mathbb{P}\,\middle\|\,\Pi\mathbb{P}^{\mathrm{R}}\right).$ (18)
Then by the data-processing inequality,
$\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\;\geq\;D_{\mathrm{KL}}\!\left(\Pi\mathbb{P}\,\middle\|\,\Pi\mathbb{P}^{\mathrm{R}}\right)=\Sigma_{\mathrm{obs}}$ (19)
holds [12, 13].
Next consider a joint system $(X,Y)$. Let $\mathbb{P}_{XY},\mathbb{P}_{XY}^{\mathrm{R}}$ be the forward and reverse path measures of the joint system, and let $\mathbb{P}_{X},\mathbb{P}_{X}^{\mathrm{R}}$ be those of the subsystem $X$. Define the joint and marginal dissipations by
$\Sigma_{XY}$ $:=D_{\mathrm{KL}}\!\left(\mathbb{P}_{XY}\,\middle\|\,\mathbb{P}_{XY}^{\mathrm{R}}\right),$ (20)
$\Sigma_{X}$ $:=D_{\mathrm{KL}}\!\left(\mathbb{P}_{X}\,\middle\|\,\mathbb{P}_{X}^{\mathrm{R}}\right).$ (21)
Their difference
$\Sigma_{\mathrm{hidden}}^{(X)}:=\Sigma_{XY}-\Sigma_{X}$ (22)
is called the “hidden dissipation from the viewpoint of XX.”
Proposition 3 (Projection-gap representation and nonnegativity of hidden dissipation).
Assume the joint forward and reverse path measures $\mathbb{P}_{XY},\mathbb{P}_{XY}^{\mathrm{R}}$ admit densities with respect to a common reference measure, and conditional path measures $\mathbb{P}_{Y|X},\mathbb{P}_{Y|X}^{\mathrm{R}}$ exist. Then the hidden dissipation $\Sigma_{\mathrm{hidden}}^{(X)}$ admits the representation
$\Sigma_{\mathrm{hidden}}^{(X)}=\int D_{\mathrm{KL}}\!\left(\mathbb{P}_{Y|X=\gamma}\,\middle\|\,\mathbb{P}_{Y|X=\gamma}^{\mathrm{R}}\right)\,\mathbb{P}_{X}(\,\mathrm{d}\gamma)\;\geq\;0.$ (23)
That is, the difference between joint and marginal dissipation is the expectation of the KL divergence between conditional path measures, and is always nonnegative.
Proof.
Write the joint path measures as $\mathrm{d}\mathbb{P}_{XY}=\,\mathrm{d}\mathbb{P}_{X}\,\,\mathrm{d}\mathbb{P}_{Y|X}$ and $\mathrm{d}\mathbb{P}_{XY}^{\mathrm{R}}=\,\mathrm{d}\mathbb{P}_{X}^{\mathrm{R}}\,\,\mathrm{d}\mathbb{P}_{Y|X}^{\mathrm{R}}$. By the chain rule for KL divergence,
$\Sigma_{XY}$ $=\int\log\!\left(\frac{\,\mathrm{d}\mathbb{P}_{X}}{\,\mathrm{d}\mathbb{P}_{X}^{\mathrm{R}}}\right)\,\mathrm{d}\mathbb{P}_{X}+\int\left[\int\log\!\left(\frac{\,\mathrm{d}\mathbb{P}_{Y|X}}{\,\mathrm{d}\mathbb{P}_{Y|X}^{\mathrm{R}}}\right)\,\mathrm{d}\mathbb{P}_{Y|X}\right]\,\mathrm{d}\mathbb{P}_{X}$
$=\Sigma_{X}+\int D_{\mathrm{KL}}\!\left(\mathbb{P}_{Y|X=\gamma}\,\middle\|\,\mathbb{P}_{Y|X=\gamma}^{\mathrm{R}}\right)\,\mathbb{P}_{X}(\,\mathrm{d}\gamma).$ (24)
The integrand is a KL divergence and hence nonnegative; integrating it against $\mathbb{P}_{X}$ preserves nonnegativity, proving (23). ∎
This proposition clarifies that information terms and mutual-information flow appearing in Ito–Sagawa and Horowitz–Esposito-type frameworks can be interpreted as concrete closed-form evaluations of $\Sigma_{\mathrm{hidden}}^{(X)}$ in specific model classes [14, 15, 16].
4.3 Mutual information and correlation free energy
At the static level, for a joint distribution $p_{XY}$ with marginals $p_{X},p_{Y}$, the mutual information
$I(X;Y):=D_{\mathrm{KL}}\!\left(p_{XY}\,\middle\|\,p_{X}p_{Y}\right)$ (25)
is the KL gap induced by the “independence projection” $p_{XY}\mapsto p_{X}p_{Y}$.
Assume the Hamiltonian is additive,
$E_{XY}(x,y)=E_{X}(x)+E_{Y}(y),$ (26)
and let $q_{X},q_{Y}$ be canonical distributions at temperature $T$:
$q_{X}(x)\propto e^{-\beta E_{X}(x)},\qquad q_{Y}(y)\propto e^{-\beta E_{Y}(y)},\qquad\beta=1/T.$ (27)
Define the nonequilibrium free-energy functionals
$F_{X}[p_{X}]:=\sum_{x}p_{X}(x)E_{X}(x)-TS(p_{X}),\qquad F_{Y}[p_{Y}]:=\sum_{y}p_{Y}(y)E_{Y}(y)-TS(p_{Y}),$ (28)
$F_{XY}[p_{XY}]:=\sum_{x,y}p_{XY}(x,y)\bigl(E_{X}(x)+E_{Y}(y)\bigr)-TS(p_{XY}),$ (29)
where $S(\cdot)$ is the Shannon entropy.
Proposition 4 (Decomposition of correlation free energy).
Under the above setting, the joint nonequilibrium free energy decomposes as
$F_{XY}[p_{XY}]=F_{X}[p_{X}]+F_{Y}[p_{Y}]+TI(X;Y).$ (30)
Moreover, the KL divergence to the product reference $q_{X}q_{Y}$ decomposes as
$D_{\mathrm{KL}}\!\bigl(p_{XY}\,\big\|\,q_{X}q_{Y}\bigr)=I(X;Y)+D_{\mathrm{KL}}(p_{X}\|q_{X})+D_{\mathrm{KL}}(p_{Y}\|q_{Y}).$ (31)
Hence the “total joint nonequilibrium free energy” $TD_{\mathrm{KL}}(p_{XY}\|q_{X}q_{Y})$ decomposes exactly into subsystem nonequilibrium free energies and the correlation free energy $TI(X;Y)$.
Proof.
Using the entropy identity $S(p_{XY})=S(p_{X})+S(p_{Y})-I(X;Y)$,
$F_{XY}[p_{XY}]$ $=\sum_{x}p_{X}(x)E_{X}(x)+\sum_{y}p_{Y}(y)E_{Y}(y)-T\{S(p_{X})+S(p_{Y})-I(X;Y)\}$
$=F_{X}[p_{X}]+F_{Y}[p_{Y}]+TI(X;Y),$ (32)
proving (30).
Also,
$D_{\mathrm{KL}}\!\bigl(p_{XY}\,\big\|\,q_{X}q_{Y}\bigr)=\sum_{x,y}p_{XY}(x,y)\log\frac{p_{XY}(x,y)}{q_{X}(x)q_{Y}(y)}.$
Writing $p_{XY}(x,y)=p_{X}(x)p_{Y|X}(y|x)$ and rearranging yields
$D_{\mathrm{KL}}\!\bigl(p_{XY}\,\big\|\,q_{X}q_{Y}\bigr)=I(X;Y)+D_{\mathrm{KL}}(p_{X}\|q_{X})+D_{\mathrm{KL}}(p_{Y}\|q_{Y}),$
which is the static counterpart of the KL chain rule. ∎
Our framework extends beyond canonical references $q_{X},q_{Y}$ to general priors $\tau$ and Min-KL solutions with information-volume constraints, so (30) and (31) admit interpretations at a more abstract level involving measurement maps and priors.
4.4 Example: a two-state Markov-chain toy model
We consider the simplest dissipative process: a two-state Markov chain, and explicitly evaluate $\Sigma_{0,T},\Sigma_{\mathrm{obs}},\Sigma_{\mathrm{hidden}}^{(X)},\Delta S_{\mathrm{sys}},\Delta S_{\mathrm{env}},Q$.
Let the state space be $\{0,1\}$ and the transition matrix be
$P=\begin{pmatrix}1-a&a\\ b&1-b\end{pmatrix},\qquad 0<a,b<1.$ (33)
The stationary distribution is
$\pi_{0}=\frac{b}{a+b},\qquad\pi_{1}=\frac{a}{a+b}.$
We fix the initial distribution to $\pi$.
A path on $[0,1]$ is $\omega=(x_{0},x_{1})$. The forward path measure is
$\mathbb{P}(x_{0}=i,x_{1}=j)=\pi_{i}P_{ij},$
and stationarity implies the time-reversed path measure
$\mathbb{P}^{\mathrm{R}}(x_{0}=i,x_{1}=j)=\pi_{j}P_{ji}.$
The one-step dissipation is then
$\Sigma_{0,1}$ $=D_{\mathrm{KL}}\!\bigl(\mathbb{P}\,\big\|\,\mathbb{P}^{\mathrm{R}}\bigr)$
$=\sum_{i,j\in\{0,1\}}\pi_{i}P_{ij}\log\!\frac{\pi_{i}P_{ij}}{\pi_{j}P_{ji}}.$ (34)
This coincides with the standard expression for housekeeping dissipation in a two-state system.
We compare two extreme choices of the measurement map.
(a) Full observation.
Take $M_{\mathrm{id}}(x)=x$. Then $\Pi\mathbb{P}=\mathbb{P}$ and $\Pi\mathbb{P}^{\mathrm{R}}=\mathbb{P}^{\mathrm{R}}$, so
$\Sigma_{\mathrm{obs}}=D_{\mathrm{KL}}(\mathbb{P}\|\mathbb{P}^{\mathrm{R}})=\Sigma_{0,1},$
and $\Sigma_{\mathrm{hidden}}^{(X)}=0$.
If $\tau$ is chosen as the uniform measure on $\{0,1\}$, then $S_{M_{\mathrm{id}},\tau}$ coincides with Shannon entropy. Since both the initial and final distributions equal $\pi$,
$\Delta S_{\mathrm{sys}}=0,\qquad\Delta S_{\mathrm{env}}=\Sigma_{0,1}.$
Thus the representation theorem in Sec. 5 gives
$Q=T\,\Delta S_{\mathrm{env}}=T\,\Sigma_{0,1},$
so all dissipation is interpreted as environment entropy increase (heat release).
(b) Extreme coarse-graining.
Let $M_{\mathrm{triv}}$ map both states to the same label:
$M_{\mathrm{triv}}(0)=M_{\mathrm{triv}}(1)=y_{0}.$
Then the observed path is always $(y_{0},y_{0})$, and $\Pi\mathbb{P}=\Pi\mathbb{P}^{\mathrm{R}}$, hence
$\Sigma_{\mathrm{obs}}=0,\qquad\Sigma_{\mathrm{hidden}}^{(X)}=\Sigma_{0,1}.$
Moreover, the observational entropy is identically zero (taking $\tau(\Omega)=1$), and
$\Delta S_{\mathrm{sys}}=0,\qquad\Delta S_{\mathrm{env}}=\Sigma_{0,1}.$
This example shows, in the simplest setting, that while $\Sigma_{0,1}$ is invariant, the values of observational dissipation $\Sigma_{\mathrm{obs}}$ and hidden dissipation $\Sigma_{\mathrm{hidden}}^{(X)}$ can vary drastically with the choice of observation $M$. In other words, $\Sigma_{0,T}$ is gauge-invariant, and choosing $(M,\tau)$ corresponds to a bookkeeping freedom that allocates dissipation between system and environment.
5 An information-theoretic definition of “heat”: a representation theorem from dissipation decomposition
So far we have (i) defined dissipation as the path-space KL divergence $\Sigma_{0,T}=D_{\mathrm{KL}}(\mathbb{P}\|\mathbb{P}^{\mathrm{R}})$ (Eq. (16)), and (ii) defined system entropy statically via observational entropy $S_{M,\tau}(p)$. In this section we reconstruct the conventional heat-like quantity $Q$, often introduced as “energy exchange with a bath,” as a representation theorem of dissipation decomposition that does not rely on LDB.
Define the system entropy change between times 0 and $T$ by
$\Delta S_{\mathrm{sys}}\;:=\;S_{M,\tau}(p_{T})-S_{M,\tau}(p_{0}).$ (35)
The dissipation on the same interval is, by definition,
$\Sigma_{0,T}\;:=\;D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\;\geq 0.$ (36)
Theorem 5 (Environment entropy change and a representation theorem for heat).
Given a measurement map $M$, a reference measure $\tau$, and path measures $\mathbb{P},\mathbb{P}^{\mathrm{R}}$, define $\Delta S_{\mathrm{sys}}$ by (35) and $\Sigma_{0,T}$ by (36). Let
$\Delta S_{\mathrm{env}}:=\Sigma_{0,T}-\Delta S_{\mathrm{sys}}.$ (37)
Then the identity
$\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\Delta S_{\mathrm{env}}$ (38)
holds tautologically. Moreover, if temperature $T$ is defined as the static Min-KL dual variable via $1/T=\beta$, we may define
$Q:=T\,\Delta S_{\mathrm{env}}.$ (39)
In this way the conventional heat quantity is recovered as
$Q=T\bigl(\Sigma_{0,T}-\Delta S_{\mathrm{sys}}\bigr),$ (40)
without invoking model-dependent assumptions such as local detailed balance.
Proof.
By definition (37), the decomposition $\Sigma_{0,T}=\Delta S_{\mathrm{sys}}+\Delta S_{\mathrm{env}}$ is an identity. Since temperature $T$ is defined by $\beta=1/T$ in Sec. 3.3, multiplying (37) by $T$ and defining $Q$ by (39) yields
$Q=T\,\Sigma_{0,T}-T\,\Delta S_{\mathrm{sys}}.$
Thus $Q/T$ equals “total dissipation minus system entropy change.” ∎
Remark 6.
The theorem is an information-theoretic representation. To identify $Q$ with the actual energy exchange with a bath, one must separately specify a dynamical definition of energy and a correspondence with a bath model. The crucial point is that, even without assuming LDB or canonical bath statistics, dissipation, observational entropy, and temperature together determine a unique candidate $Q$ as a representation. In subsequent applied work, we plan to evaluate (39) for chemical reaction networks and RNAP–DNA contract-like systems, and compare with conventional definitions of heat and Landauer-type principles.
5.1 Consistency with the cross-entropy interpretation of heat
In Sec. 5, we defined total dissipation as the path relative entropy $\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)$ and adopted observational entropy $S_{M,\tau}$ as system entropy. We then introduced heat $Q$ via
$\Delta S_{\mathrm{env}}:=\Sigma_{0,T}-\Delta S_{\mathrm{sys}},\qquad Q:=T\,\Delta S_{\mathrm{env}}.$ (41)
Here we clarify, conceptually, the consistency of this definition with the frequently used viewpoint that “heat (or its dimensionless form) can be expressed as a cross-entropy.”
(i) Agreement with the standard identification in isothermal environments
In standard stochastic thermodynamics (e.g. when an isothermal bath at temperature $T$ and a mechanical definition of heat $q_{\mathrm{ex}}$ are given), dissipation is written as
$\Sigma=\Delta S_{\mathrm{sys}}+\beta q_{\mathrm{ex}},\qquad\beta=\frac{1}{T}.$ (42)
Thus
$\beta q_{\mathrm{ex}}=\Sigma-\Delta S_{\mathrm{sys}}.$ (43)
Therefore (41) implies
$\frac{Q}{T}=\Delta S_{\mathrm{env}}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}}.$ (44)
Hence $Q$ can be identified with the conventional (equivalent) heat $q_{\mathrm{ex}}$ in isothermal stochastic thermodynamics. The difference is conceptual: we define $Q$ from dissipation decomposition, and only later (if desired) impose model-dependent identifications such as (42).
(ii) Cross-entropy as a readable form
The cross-entropy between distributions $p$ and $q$ is
$H(p,q):=-\sum_{x}p(x)\log q(x),$ (45)
and satisfies
$H(p,q)=S(p)+D_{\mathrm{KL}}(p\|q).$ (46)
Thus, when a process naturally singles out a reference (target) distribution $p^{\star}$ and the dimensionless heat is written as
$\frac{Q}{T}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}},$ (47)
if the right-hand side can be evaluated/rearranged to contain a KL term $D_{\mathrm{KL}}(\cdot\|p^{\star})$, then (46) enables a rewriting of $Q/T$ in terms of a cross-entropy $H(\cdot,p^{\star})$. In this sense, (47) acts as a master bookkeeping identity that contains cross-entropy expressions as special readable forms.
Typical situations include:
  1. (a)
    Relaxation/erasure processes toward a target $p^{\star}$ in which dissipation (or its dominant part) reduces to a form like $D_{\mathrm{KL}}(p_{0}\|p^{\star})$. Then (46) produces $S(p_{0})+D_{\mathrm{KL}}(p_{0}\|p^{\star})=H(p_{0},p^{\star})$, allowing $Q/T$ to be interpreted as a cross-entropy.
  2. (b)
    Processes where the final state becomes (in a coarse-grained sense) sharply localized so that $S(p_{T})\simeq 0$ (e.g. logical erasure). Then $\Delta S_{\mathrm{sys}}$ is effectively fixed, and if $\Sigma_{0,T}$ contains a term $D_{\mathrm{KL}}(p_{0}\|p^{\star})$, $Q/T$ tends to simplify to $H(p_{0},p^{\star})$ (up to state-function terms).
(iii) Connection to Landauer-type arguments in Tsuruyama (2025)
Tsuruyama [17] starts from a decomposition of the form $\sigma=\Delta S+\beta q_{\mathrm{ex}}$ in a Landauer-type erasure setting and rearranges $\beta q_{\mathrm{ex}}$ to obtain, for a binary erasure problem,
$\beta q_{\mathrm{ex}}=-p\log p^{\ast}-(1-p)\log(1-p^{\ast}),$ (48)
which is a cross-entropy form. This is naturally interpreted as a concrete realization of
$\beta q_{\mathrm{ex}}=\Sigma-\Delta S_{\mathrm{sys}}$ (49)
via (46). Therefore, our representation (41) contains Tsuruyama’s “heat = cross-entropy” redefinition as a special case, without assuming a specific model class or erasure protocol at the level of definitions.
(iv) Implications of using observational entropy $S_{M,\tau}$
We define $\Delta S_{\mathrm{sys}}$ using observational entropy $S_{M,\tau}$ rather than Shannon entropy. This does not break consistency with cross-entropy interpretations. Instead, because $S_{M,\tau}$ explicitly includes coarse-graining cell volumes $\tau(C_{y})$, it naturally captures situations where “the logical state is fixed but microstate volume (information volume) remains.” When one rewrites $Q/T$ in cross-entropy form, the essential choice is the reference (target) distribution $p^{\star}$; in our language, this corresponds to the gauge choice of the measurement map $M$ and the reference measure $\tau$ (the descriptive convention).
6 Relation to existing theories and features of the present framework
This section summarizes how our theoretical core relates to existing information thermodynamics and highlights conceptual features and open problems.
6.1 Relation to existing information thermodynamics
Expressing dissipation as a path-space KL divergence,
$\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right),$
is established for reversible phase-space dynamics [9] and for general stochastic processes [10, 11]. The coarse-graining lower bound $\Sigma_{0,T}\geq\Sigma_{\mathrm{obs}}$ (Eq. (19)) appears in discussions of dissipation under coarse graining [12, 13].
On the static side, observational entropy $S_{M,\tau}(p)$—including the cell volume $\tau(C_{y})$ in the entropy definition—has been systematically developed by Safránek et al. in classical and quantum settings [3, 4, 5, 6, 7]. Our contribution is to connect observational entropy to path-KL dissipation and to dual variables in Min-KL inference, thereby unifying static and dynamic information thermodynamics under a single reference measure $\tau$.
Finally, the idea that correlations between a system and memory (or environment) store free energy quantified by $k_{\mathrm{B}}T$ times mutual information $I(X;Y)$ has been emphasized repeatedly in the literature by Sagawa–Ueda and Horowitz–Esposito and others [14, 15, 16]. We embed these discussions into a KL-geometric decomposition of $D_{\mathrm{KL}}(p_{XY}\|q_{X}q_{Y})$ (Eq. (31)), treating local and correlation free energies within a single information-geometric framework.
6.2 Conceptual features
Our main conceptual contributions can be summarized as follows.
  1. (i)
    Primary definition of pressure via information-volume constraints. We introduce volume not as geometric space volume but via the cell volume $v(y)=\tau(C_{y})$ determined by $(M,\tau)$ and the associated log-volume $\ell_{V}(x)$. Incorporating the constraint $\mathbb{E}[\ell_{V}]=\bar{v}$ into Min-KL inference defines pressure $P$ as the shadow price (dual variable), on equal footing with temperature $T$ and chemical potential $\mu$ (Eq. (10)). This repositions pressure as the variable conjugate to an observation- and convention-dependent information volume, enabling unified extensions to nonstandard state spaces and coarse-grainings.
  2. (ii)
    Hidden dissipation and information terms as projection gaps. At the path level, the difference between joint dissipation $\Sigma_{XY}$ and marginal dissipation $\Sigma_{X}$ is expressed as an expectation of a KL divergence between conditional path measures (Eq. (23)). Thus hidden dissipation $\Sigma_{\mathrm{hidden}}^{(X)}$ is not an ad hoc correction term but an intrinsic KL gap induced by projection from the joint description to a marginal one. Information terms and mutual-information flows in existing information thermodynamics can be understood as special closed-form evaluations of this projection gap.
  3. (iii)
    Heat reconstructed as an information-theoretic representation theorem. With dissipation defined by path KL, system entropy defined by observational entropy, and temperature defined as a Min-KL dual variable, we define $\Delta S_{\mathrm{env}}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}}$ and represent heat as
    $Q=T\bigl(\Sigma_{0,T}-\Delta S_{\mathrm{sys}}\bigr).$
    In this viewpoint, $Q$ is not a primitive notion but a secondary representation derived from dissipation decomposition and dual variables, disentangled from model-dependent descriptions such as “energy exchange with a bath.”
6.3 Open problems and directions
We conclude by listing several theoretical and applied directions.
  1. (a)
    Phase-transition structure with information-volume constraints. In Min-KL inference with $\mathbb{E}[\ell_{V}]=\bar{v}$, it is important to analyze where the dual map $(\beta,\alpha,\pi)\mapsto(U,\bar{N},\bar{v})$ loses regularity and when ensemble nonequivalence occurs, which directly relates to phase transitions involving information volume. In particular, effective models without geometric volume or with nonlocal measurements $M$ may display novel critical phenomena.
  2. (b)
    Optimizing observational dissipation bounds and observation design. The gap between observational dissipation $\Sigma_{\mathrm{obs}}$ and true dissipation $\Sigma_{0,T}$ (and hence hidden dissipation) is sensitive to $(M,\tau)$. Designing $(M,\tau)$ to maximize $\Sigma_{\mathrm{obs}}$ under experimental constraints is a natural optimization problem linking statistical efficiency of dissipation estimation to experimental design.
  3. (c)
    New decomposition formulas combining the first-law-type relation and path dissipation. Combining the static first-law-type relation (15) with path-KL dissipation $\Sigma_{0,T}$ suggests new decompositions/inequalities that simultaneously involve changes in internal energy, information volume, and particle number together with dissipation, correlation generation, and hidden dissipation. For instance, deriving second-law-type inequalities including an information-volume work-like term and the production of correlation free energy $TI(X;Y)$ under a general prior $\tau$ would further tighten the link between nonequilibrium thermodynamics and information theory.
  4. (d)
    Applications to concrete models and comparisons with existing principles. In planned applications to RNAP–DNA systems and chemical reaction systems analyzed via large deviation theory [17], explicitly evaluating correlation free energy, hidden dissipation, and the heat representation (39) is essential for comparing with fluctuation theorems and Landauer-type principles. Clarifying when conventional heat/work identifications agree with (or differ from) our information-volume/dissipation-based identifications is also crucial for interpretation in applications.
In summary, our information-volume constraint approach encompasses existing results in information thermodynamics [9, 10, 11, 12, 13, 3, 4, 5, 6, 7, 14, 15, 16, 17] while explicitly elevating $(M,\tau)$ to primitive building blocks. This repositions static and dynamic information thermodynamics within a single KL-geometric framework and provides an abstract foundation for future applications and experimental connections.